How many hydrogen bonds exist between this DNA strand and its complementary strand?

5′–GCTCACG–3′

number of hydrogen bonds between strands: ?
Answer #1

According to Watson and Crick model of base pairs, adenine and thymine have 2 hydrogen bonds in between them, where as guanine and cytosine have 3 hydrogen bonds in between them. The given DNA strand has one adenine and one thymine, both of them will interact with the complementary strand of the DNA by two hydrogen bonds with their complementary base pair. The rest 5 bases, i.e, 3*cytosine and 2*guanine both interact with their complementary base pairs in the complementary strand by 3 hydrogen bonds each.

Total hydrogen bonds = 2*2 + 5*3 =4 +15 =19

So the DNA strand given will form 19 hydrogen bonds with the complementary DNA strand according to Watson Crick model of base pairs.